OpenStudy (anonymous):

If a coin is flipped 10 times, what is the probability that it will show all heads or all tails?

OpenStudy (perl):

so you want the probability of HHH...H and TTT...T each flip has probability of 1/2 , and we multiply probabilities P(H & H & H ...) = P( H) * P(H) * .. P(H) = 1/2 * 1/2 * ... 1/2

OpenStudy (anonymous):

i beleive that it is .000976563% but i could be wrong

OpenStudy (perl):

P ( HHH HHH HHH H TTT TTT TTT T) = (1/2)^10 + (1/2)^10

OpenStudy (perl):

P ( HHHHHHHHHH or TTTTTTTTTT) = (1/2)^10 + (1/2)^10 = 2 * (1/2)^10 = 1/2^9

OpenStudy (perl):

you are wrong, it should be .001953

OpenStudy (anonymous):

hmm sorry its been awhile as i am in the 11th grade

OpenStudy (anonymous):

a 5:5 chance since half of 10 is 5

OpenStudy (perl):

there is a 5:5 chance with 1 coin getting heads, not ten coins

OpenStudy (perl):

think about it, it is extremely unlikely to flip a coin ten times and get ten consecutive heads, or ten consecutive tails